OpenStudy (anonymous):

Determine u*v if ||u||=5, ||v||=4, and theta=pi/3, where theta is the angle between u and v.

OpenStudy (anonymous):

<u,v>=||u||.||v||.cos(theta)

OpenStudy (savvy):

u*v = lullvlcospi/3 = 5.4.1/2 =10

OpenStudy (anonymous):

my choices are: 14.14 13.66 10.00 12.07 17.32

OpenStudy (anonymous):

so <u,v>=5.4.cos(pi/3)

OpenStudy (anonymous):

10

OpenStudy (anonymous):

when i plugged in 5.4cos(pi/3), i got 5.399

jimthompson5910 (jim_thompson5910):

remember to convert to radian mode

OpenStudy (anonymous):

you put your calc in radian mode?

OpenStudy (anonymous):

well now i got 2.7

jimthompson5910 (jim_thompson5910):

and it's not 5.4 it's 5*4

OpenStudy (anonymous):

ohhhh

jimthompson5910 (jim_thompson5910):

yeah dots and multiplication are confused a lot

OpenStudy (anonymous):

i got 10 now, thanks for the help!

jimthompson5910 (jim_thompson5910):

sure thing